OpenStudy (anonymous):

Can someone help me simplify this 1 radical? Thanks!

jimthompson5910 (jim_thompson5910):

Are you familiar with perfect cube numbers?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

If you could provide an answer, with how you got it, so I can study that, it would be helpful!

jimthompson5910 (jim_thompson5910):

you're familiar with perfect squares right? 1, 4, 9, 16, 25, ... numbers like that

jimthompson5910 (jim_thompson5910):

Well anyway, the list of perfect cubes are 1, 8, 27, 64, 125, ... since 1^3 = 1 2^3 = 8 3^3 = 27 4^3 = 64 5^3 = 125 .... ....

jimthompson5910 (jim_thompson5910):

you then use this list to help you simplify each radical since 54 = 27*2, we can say this \[\large \sqrt[3]{54} = \sqrt[3]{27*2}\] \[\large \sqrt[3]{54} = \sqrt[3]{27}*\sqrt[3]{2}\] \[\large \sqrt[3]{54} = \sqrt[3]{3^3}*\sqrt[3]{2}\] \[\large \sqrt[3]{54} = 3*\sqrt[3]{2}\]

jimthompson5910 (jim_thompson5910):

You apply the same basic steps to the other radicals. Once they are simplified fully, you combine like terms